Grade Select GradeDifferential Calculus

​Q. FIND THE EQUATION OF THE NORMAL TO THE CURVE-
x3 + y3 = 8xy
AT THE POINT WHERE IT MEETS THE CURVE y2 = 4x . OTHER THAN THE ORIGIN.

Bharat Makkar
11 Years agoGrade Select Grade

1 Answer

Jitender Singh
11 Years ago
Ans:
Hello Student,
Please find answer to your question below

Lets find the point of intersection:
x^3 + y^3 = 8xy…......(1)
x^3 = y(8x-y^2)
y^2 = 4x
x^3 = y(8x-4x)
x^3 = 4xy
x^2 = 4y
Put in (1)
x^3 + (\frac{x^2}{4})^3 = 8x.\frac{x^2}{4}
x^3 + \frac{x^6}{64} = 2x^3
\frac{x^6}{64} = x^3
x^3(x^3-64) = 0
x = 0, 4

Point of intersection other than origin
(4, 4)
Lets find the slope of normal:
\frac{-1}{y'}
Differentiate (1)
3x^2 + 3y^2y' = 8y+8xy'
3x^2 - 8y = (8x-3y^2)y'
3.4^2 - 8.4 = (8.4-3.4^2)y'
16 = (-16)y'
y' = -1
Slope: 1
Equation of normal:
(y-4)=(x-4)
y=x